OpenStudy (anonymous):

Need help with 5/8 + 1/6

OpenStudy (anonymous):

you'll need to find a common denominator so you can add the fractions

OpenStudy (anonymous):

the easiest way to do that is by multiplying the denominators to find it,

OpenStudy (anonymous):

The LCM of 8 and 6 is 24

OpenStudy (anonymous):

We adjust the fractions so they have a denominator of 24, and we also have to change the numerator

OpenStudy (anonymous):

5/8 = 15/24, and 1/6 = 4/24. So what is the answer to 15/24 + 4/24?

OpenStudy (anonymous):

Since they have the same denominator, you just add the numerators.

OpenStudy (anonymous):

19/24 is my answer also Thanks!

OpenStudy (anonymous):

Good job.

OpenStudy (anonymous):

awesome team work you guys @AnswerMyQuestions @iriscar3

OpenStudy (anonymous):

Thanks